Hi I am bnot sure how to do these two problems. I hope some one can help me. Thank you very much.
1. Solve by factoring:
x2 – 9x = -8
2. Solve using the square root property:
2x2 – 5 = 93

Answer by checkley77(12844)

x2–9x=-8
X^2-9X+8=0
(X-1)(X-8)=0
X-1=0
X=1 ANSWER.
X-8=0
X-8 ANSWER.
---------------------------
2x2–5=93
2X^2=93+5
2X^2=98
X^2=98/2
X^2=49
X=SQRT49
X=7 ANSWER.